OpenStudy (anonymous):

Find the value of 7x^2 / sqrt x^5 7 sqrt x / x 7 /x 7 sqrt x 7x / sqrt x

OpenStudy (tyteen4a03):

\(\frac{7x^2}{\sqrt{x^5}}\) Remember that \(\sqrt{x^y}\) = x^(y/2)

OpenStudy (anonymous):

ok but im still puzzeled

OpenStudy (anonymous):

sqrt of x^5 is x^2*sqrtx. So, this gives 7/sqrtx

OpenStudy (anonymous):

\[\Large{7x^2\over x^{5\over2}}=7x^2\times x^{-5\over2}=7\times x^{2-{5\over2}}\]

OpenStudy (tyteen4a03):

@electrokid Your LaTeX is better than me. :P

OpenStudy (anonymous):

\[=\Large{7\times x^{4-5\over2}}=7\times x^{-1\over2}={7\over x^{1\over2}}\]

OpenStudy (anonymous):

@tyteen4a03 :D. That is what I use all the time. Got rid of "Office suites" a long time ago